### v1

# Hadronic molecules with four charm or beauty quarks

Abstract
We apply the extended local hidden gauge formalism to study the meson-meson interactions with the quark constituents $cc\bar{c}\bar{c}$, $cc\bar{c}\bar{b}/\bar{c}\bar{c}cb$, $cc\bar{b}\bar{b}/\bar{c}\bar{c}bb$, $bb\bar{c}\bar{b}/\bar{b}\bar{b}cb$, and $bb\bar{b}\bar{b}$, where the exchanged mesons can only be the fully-heavy vector mesons $J/\psi$, $B_{c}^{*}$, and $\Upsilon$. We solve the coupled-channel Bethe-Salpeter equation to derive two poles in the $bb\bar{c}\bar{b}$ system and two poles in the $cc\bar{c}\bar{b}$ system (with four charge-conjugated poles in the $\bar{b}\bar{b}cb$ and $\bar{c}\bar{c}cb$ systems): one pole existing in the $bb\bar{c}\bar{b}$ system corresponds to the sub-threshold bound state when setting the cutoff momentum Λ>850 $\Lambda>850\textrm{ MeV}$, while it causes the threshold effect with Λ<850 $\Lambda<850\textrm{ MeV}$; the other pole existing in this system corresponds to the sub-threshold bound state with Λ>1100 $\Lambda>1100\textrm{ MeV}$; the two poles existing in the $cc\bar{c}\bar{b}$ system correspond to the sub-threshold bound states only when setting Λ>1550 $\Lambda>1550\textrm{ MeV}$ and Λ>2650 $\Lambda>2650\textrm{ MeV}$, making them difficult to be identified as deeply-bound hadronic molecules. We propose to investigate the two poles of the $bb\bar{c}\bar{b}$ system in the $\mu^{+}\mu^{-}B_{c}^{-}$ channel at LHC.
fully-heavy hadronic molecule, Bethe-Salpeter equation, local hidden-gauge formalism
I Introduction
In recent decades, the study of exotic hadrons has gradually become a focal point in hadron physics. Some exotic hadrons exhibit multiquark compositions, such as the compact tetraquark states and the meson-meson molecular states. The picture of hadronic molecules has achieved significant success in the light-quark sector Oller:1997ti ; Oller:1998hw ; Oset:1997it ; Jido:2003cb ; Bruns:2010sv ; Garcia-Recio:2003ejq ; Hyodo:2002pk , which can be used to explain many resonances, such as the $f_{0}(980)$ and $a_{0}(980)$, etc. Besides, many hidden-charm pentaquark states observed in the past decade can be interpreted as the hadronic molecules that are dynamically generated through the meson-baryon interactions within the local hidden gauge framework Wu:2010jy ; Wu:2010vk ; Chen:2015sxa ; He:2015cea ; Xiao:2013yca ; Roca:2015dva ; Liu:2015fea ; Uchino:2015uha . In recent years, several exotic structures in the di-$J/\psi$ invariant mass spectrum have been reported by the LHCb, CMS, and ATLAS collaborations LHCb:2020bwg ; ATLAS:2023bft ; CMS:2023owd , including the $X(6200)$, $X(6600)$, $X(6900)$, and $X(7200)$. These structures are good candidates for the fully-charmed tetraquark states. Extensive theoretical investigations have been performed to elucidate their nature liu:2020eha ; Tiwari:2021tmz ; Lu:2020cns ; Faustov:2020qfm ; Zhang:2020xtb ; Li:2021ygk ; Bedolla:2019zwg ; Weng:2020jao ; Liu:2021rtn ; Giron:2020wpx ; Karliner:2020dta ; Zhao:2020zjh ; Mutuk:2021hmi ; Wang:2021kfv ; Wang:2020ols ; Ke:2021iyh ; Zhu:2020xni ; Jin:2020jfc ; Yang:2021hrb ; Albuquerque:2020hio ; Albuquerque:2021erv ; Wu:2022qwd ; Asadi:2021ids ; Yang:2020wkh ; Feng:2020riv ; Ma:2020kwb ; Maciula:2020wri ; Goncalves:2021ytq ; Wang:2020gmd ; Esposito:2021ptx ; Zhuang:2021pci ; Zhao:2020nwy ; Becchi:2020uvq ; Sonnenschein:2020nwn ; Zhu:2020snb ; Wan:2020fsk ; Gordillo:2020sgc ; Liu:2020tqy ; Majarshin:2021hex ; Kuang:2022vdy ; Wang:2021mma ; Chen:2016jxd ; Czarnecki:2017vco , some of which attempted to explore their nature as molecular states Guo:2020pvt ; Cao:2020gul ; Gong:2020bmg ; Dong:2021lkh ; Ortega:2023pmr ; Wang:2022yes ; Zhou:2022xpd , but a definitive and conclusive resolution remains elusive.
Previous theoretical studies on the fully-heavy tetraquark states mainly focus on the interpretation of compact tetraquark states, while there are not so many studies based on the interpretation of hadronic molecular states. This is because the exchanged hadrons of these systems all have quite large masses, such as the fully-heavy vector mesons $J/\psi$, $B_{c}^{*}$, and $\Upsilon$ exchanged in the $cb\bar{c}\bar{b}$ system within the extended local hidden gauge formalism, so their induced interactions are significantly suppressed. In Ref. Liu:2023gla we studied the $cb\bar{c}\bar{b}$ system to explore the existence of the fully-heavy hadronic molecules $B_{c}^{(*)}\bar{B}_{c}^{(*)}$. Within the extended local hidden gauge framework, we found that the two fully-heavy mesons $B_{c}^{(*)}$ and $\bar{B}_{c}^{(*)}$ are able to form a bound state by exchanging the relatively lighter meson $J/\psi$.
In this paper we apply the extended local hidden gauge formalism to further investigate the fully-heavy hadronic molecules existing in the $cc\bar{c}\bar{c}$, $cc\bar{c}\bar{b}/\bar{c}\bar{c}cb$, $cc\bar{b}\bar{b}/\bar{c}\bar{c}bb$, $bb\bar{c}\bar{b}/\bar{b}\bar{b}cb$, and $bb\bar{b}\bar{b}$ systems. By solving the coupled-channel Bethe-Salpeter equation, we evaluate the hadronic molecules generated from the meson-meson interactions in these systems. Our results suggest the existence of two potential bound states in the $bb\bar{c}\bar{b}$ system (with two charge-conjugated states in the $\bar{b}\bar{b}cb$ system), while they manifest as the threshold effects as the cutoff momentum varies. Both of these two structures share the same spin-parity quantum number $J^{P}=1^{+}$, and their possible observing channel is $\mu^{+}\mu^{-}B_{c}^{-}$ at LHC. Besides, we find two poles in the $cc\bar{c}\bar{b}$ system (with two charge-conjugated poles in the $\bar{c}\bar{c}cb$ system), but it is difficult to identify them as deeply-bound hadronic molecules.
This paper is organized as follows. In Sec. II we apply the local hidden gauge formalism to derive the potentials for the interactions of the $cc\bar{c}\bar{c}$, $cc\bar{c}\bar{b}/\bar{c}\bar{c}cb$, $cc\bar{b}\bar{b}/\bar{c}\bar{c}bb$, $bb\bar{c}\bar{b}/\bar{b}\bar{b}cb$, and $bb\bar{b}\bar{b}$ systems. Based on the obtained results, we solve the coupled-channel Bethe-Salpeter equation in Sec. III to extract the poles, some of which may qualify as fully-heavy hadronic molecules. A brief summary is given in Sec. IV.
II formalism
In Ref. Liu:2023gla we have applied the extended local hidden gauge formalism to study the interactions of the $cb\bar{c}\bar{b}$ system. In this section we follow the same approach to study the $cc\bar{c}\bar{c}$, $cc\bar{c}\bar{b}/\bar{c}\bar{c}cb$, $cc\bar{b}\bar{b}/\bar{c}\bar{c}bb$, $bb\bar{c}\bar{b}/\bar{b}\bar{b}cb$, and $bb\bar{b}\bar{b}$ systems. Within this framework, the interactions are primarily contributed by the vector meson exchange, as depicted in Fig. 1. The corresponding Lagrangians are written as:
$\mathcal{L}_{VPP}$ $=$ $-ig\,\langle[P,\partial_{\mu}P]V^{\mu}\rangle\,,$
$\mathcal{L}_{VVV}$ $=$ $ig\,\langle(V^{\mu}\partial_{\nu}V_{\mu}-\partial_{\nu}V^{\mu}V_{\mu})V^{\nu}\rangle\,,$ (1)
$\mathcal{L}_{VVVV}$ $=$ $\frac{g^{2}}{2}\langle V_{\mu}V_{\nu}V^{\mu}V^{\nu}-V_{\nu}V_{\mu}V^{\mu}V^{\nu}\rangle\,,$
where
$P=\left(\begin{array}[]{cc}\eta_{c}&B_{c}^{+}\\ B_{c}^{-}&\eta_{b}\end{array}\right),~{}~{}~{}V=\left(\begin{array}[]{cc}J/\psi&B_{c}^{*+}\\ B_{c}^{*-}&\Upsilon\end{array}\right)\,.$ (2)
The coupling constant $g$ is generally defined as $g={M_{V}}/({2f_{P}})$, with $M_{V}$ the mass of the exchanged vector meson and $f_{P}$ the decay constant of its corresponding pseudoscalar meson. Since the charm and bottom quarks do not form a flavor $SU(2)$ symmetry, we can not use an overall parameter any more. For the exchange of the $J/\psi$, $B_{c}^{*}$, and $\Upsilon$ mesons, we respectively use
[…]
Besides, we use $g^{4}=g_{V_{1}}g_{V_{2}}g_{V_{3}}g_{V_{4}}$ for the contact term, with $V_{1\cdots 4}$ the four connected vector mesons.
Figure 1: The meson-meson interactions arising from the vector meson exchange: (a) between two pseudoscalar mesons, (b) between one vector meson and one pseudoscalar meson, and (c) between two vector mesons. The subfigure (d) describes the contact term connecting four vector mesons.
We derive the following interaction potentials from the Lagrangians given in Eq. (1) as,
$V_{PP}(s)$ $=$ $C_{PP}^{t}\times g^{2}(p_{1}+p_{3})(p_{2}+p_{4})$
$+$ $C_{PP}^{u}\times g^{2}(p_{1}+p_{4})(p_{2}+p_{3})\,,$
$V_{VP}(s)$ $=$ $C_{VP}^{t}\times g^{2}(p_{1}+p_{3})(p_{2}+p_{4})~{}\epsilon_{1}\cdot\epsilon_{3}$
$+$ $C_{VP}^{u}\times g^{2}(p_{1}+p_{4})(p_{2}+p_{3})~{}\epsilon_{1}\cdot\epsilon_{3}\,,$
$V_{VV}(s)$ $=$ $V^{ex}_{VV}(s)+V^{co}_{VV}(s)\,,$ (6)
$V^{ex}_{VV}(s)$ $=$ $C_{VV}^{t}\times g^{2}(p_{1}+p_{3})(p_{2}+p_{4})\epsilon_{1}\cdot\epsilon_{3}\epsilon_{2}\cdot\epsilon_{4}$
$+$ $C_{VV}^{u}\times g^{2}(p_{1}+p_{3})(p_{2}+p_{4})\epsilon_{1}\cdot\epsilon_{4}\epsilon_{2}\cdot\epsilon_{3}\,,$
where $p_{1}$ and $p_{2}$ are four-momenta of the incoming mesons; $p_{3}$ and $p_{4}$ are four-momenta of the outgoing mesons; $\epsilon_{1}$ and $\epsilon_{2}$ are polarization vectors of the incoming mesons; $\epsilon_{3}$ and $\epsilon_{4}$ are polarization vectors of the outgoing mesons. We use the subscripts $PP$, $VP$, and $VV$ to denote the pseudoscalar-pseudoscalar, vector-pseudoscalar, and vector-vector sectors, respectively. We use the superscripts $t$, $u$, and $co$ to denote the vector meson exchange in the $t$ and $u$ channels as well as the contact term, respectively.
We further derive the scattering amplitudes from Eqs. (II-II) by solving the Bethe-Salpeter equation as,
$T_{PP/VP/VV}=\left({\bf 1}-V_{PP/VP/VV}\bullet G\right)^{-1}\bullet V_{PP/VP/VV},$ (8)
where $G(s)$ is the diagonal loop function, whose expression for the $i$th channel is
$G_{ii}(s)=i\int{\frac{d^{4}q}{(2\pi)^{4}}\frac{1}{q^{2}-m_{1}^{2}+i\epsilon}\frac{1}{(p-q)^{2}-m_{2}^{2}+i\epsilon}}\,.$ (9)
In the above expression, $m_{1,2}$ are the masses of the two mesons involved in this channel, and $s=p^{2}$ with $p$ the total four-momentum. We apply the cutoff method to regularize it as
$G_{ii}(s)=\int_{0}^{\Lambda}\frac{d^{3}q}{(2\pi)^{3}}\frac{\omega_{1}+\omega_{2}}{2\omega_{1}\omega_{2}}\frac{1}{s-(\omega_{1}+\omega_{2})^{2}+i\epsilon}\,,$ (10)
[…]
We shall calculate the coefficient matrices $C_{PP/VP/VV}^{t/u/co}$ in the following subsections, separately for the $cc\bar{c}\bar{c}$, $cc\bar{c}\bar{b}/\bar{c}\bar{c}cb$, $cc\bar{b}\bar{b}/\bar{c}\bar{c}bb$, $bb\bar{c}\bar{b}/\bar{b}\bar{b}cb$, and $bb\bar{b}\bar{b}$ systems. Before doing this, we summarize their relevant coupled channels in Table 1.
Table 1: Coupled channels considered for the $cc\bar{c}\bar{c}$, $cc\bar{c}\bar{b}/\bar{c}\bar{c}cb$, $cc\bar{b}\bar{b}/\bar{c}\bar{c}bb$, $bb\bar{c}\bar{b}/\bar{b}\bar{b}cb$, and $bb\bar{b}\bar{b}$ systems.
Constituent $PP$ sector $VP$ sector $VV$ sector
$cc\bar{c}\bar{c}$ $\eta_{c}\eta_{c}$ $J/\psi\eta_{c}$ $J/\psi J/\psi$
$cc\bar{c}\bar{b}$ $\eta_{c}B_{c}^{+}$ $J/\psi B_{c}^{+},B_{c}^{*+}\eta_{c}$ $J/\psi B_{c}^{*+}$
$cc\bar{b}\bar{b}$ $B_{c}^{+}B_{c}^{+}$ $B_{c}^{*+}B_{c}^{+}$ $B_{c}^{*+}B_{c}^{*+}$
$bb\bar{c}\bar{b}$ $\eta_{b}B_{c}^{-}$ $\Upsilon B_{c}^{-},B_{c}^{*-}\eta_{b}$ $\Upsilon B_{c}^{*-}$
$bb\bar{b}\bar{b}$ $\eta_{b}\eta_{b}$ $\Upsilon\eta_{b}$ $\Upsilon\Upsilon$
II.1 The $cc\bar{c}\bar{c}$ system
In the $cc\bar{c}\bar{c}$ system, the $PP$ interaction involves only one channel $\eta_{c}\eta_{c}$, the $VP$ interaction involves only one channel $J/\psi\eta_{c}$, and the $VV$ interaction also involves only one channel $J/\psi J/\psi$.
Since the two vertices are both zero, $\mathcal{L}_{\eta_{c}\eta_{c}J/\psi}=0$ and $\mathcal{L}_{J/\psi J/\psi J/\psi}=0$, the interactions in this system all vanish within the extended local hidden gauge framework:
$C_{PP}^{t}$ $=$ $0\,,$
$C_{PP}^{u}$ $=$ $0\,,$
$C_{VP}^{t}$ $=$ $0\,,$ (11)
$C_{VV}^{t}$ $=$ $0\,,$
$C_{VV}^{u}$ $=$ $0\,,$
$C_{VV}^{co}$ $=$ $0\,.$
Therefore, our results do not support the existence of hadronic molecules in the $cc\bar{c}\bar{c}$ system.
II.2 The $cc\bar{c}\bar{b}$ and $\bar{c}\bar{c}cb$ systems
The results for the $cc\bar{c}\bar{b}$ and $\bar{c}\bar{c}cb$ systems are the same, so we only need to study the $cc\bar{c}\bar{b}$ system. In this system, the $PP$ interaction involves only one channel $\eta_{c}B_{c}^{+}$, whose coefficients are
$C_{PP}^{t}$ $=$ $0\,,$ (12)
$C_{PP}^{u}$ $=$ $\lambda\frac{1}{m^{2}_{B_{c}^{*}}}\,.$
The reduction factor $\lambda$ is introduced to account for the large mass difference between the initial and final mesons. Following Ref. Yu:2018yxl , numerically we use
$\lambda_{\eta_{c}B_{c}^{+}\to B_{c}^{+}\eta_{c}}\approx\frac{-m_{B_{c}^{*}}^{2}}{(m_{\eta_{c}}-m_{B_{c}})^{2}-m_{B_{c}^{*}}^{2}}=1.37.$ (13)
The coefficient $C_{PP}^{u}$ is positive, indicating the interaction due to the exchange of the $B_{c}^{*}$ meson to be repulsive, so there do not exist hadronic molecules in the $PP$ sector.
The $VP$ interaction involves two coupled channels $J/\psi B_{c}^{+}$ and $B_{c}^{*+}\eta_{c}$, whose coefficient is
$C_{VP}^{t}$ $=$ $\left(\begin{array}[]{c|cc}J=1&J/\psi B_{c}^{+}&B_{c}^{*+}\eta_{c}\\ \hline\cr J/\psi B_{c}^{+}&0&\lambda\frac{1}{m_{B_{c}^{*}}^{2}}\\ B_{c}^{*+}\eta_{c}&\lambda\frac{1}{m_{B_{c}^{*}}^{2}}&0\end{array}\right)\,.$ (17)
Diagonalizing this $2\times 2$ matrix, we obtain two decoupled channels:
$|VP^{+}\rangle$ $=$ $\frac{1}{\sqrt{2}}\left(|J/\psi B_{c}^{+}\rangle+|B_{c}^{*+}\eta_{c}\rangle\right),$ (18)
$|VP^{-}\rangle$ $=$ $\frac{1}{\sqrt{2}}\left(|J/\psi B_{c}^{+}\rangle-|B_{c}^{*+}\eta_{c}\rangle\right),$ (19)
whose coefficient is
$C_{VP}^{\prime t}$ $=$ $\left(\begin{array}[]{c|cc}J=1&VP^{+}&VP^{-}\\ \hline\cr VP^{+}&\lambda\frac{1}{m_{B_{c}^{*}}^{2}}&0\\ VP^{-}&0&-\lambda\frac{1}{m_{B_{c}^{*}}^{2}}\end{array}\right)\,.$ (23)
Hence, the interaction due to the exchange of the $B_{c}^{*}$ meson in the $VP^{-}$ channel turns out to be attractive, so there may exist a hadronic molecule of $J^{P}=1^{+}$ in the $VP$ sector.
The $VV$ interaction involves only one channel $J/\psi B_{c}^{*+}$, whose coefficients are
$C_{VV}^{t}$ $=$ $0\,,$ (24)
$C_{VV}^{u}$ $=$ $\lambda\frac{1}{m^{2}_{B_{c}^{*}}}\,,$
and the relevant contact term is
$V^{co}_{J/\psi B_{c}^{*+}\to J/\psi B_{c}^{*+}}(s)$ $=$ $\left\{\begin{array}[]{cc}-2g^{2}&~{}\textrm{for $J=0$},\\ 3g^{2}&~{}\textrm{for $J=1$},\\ g^{2}&~{}\textrm{for $J=2$}.\end{array}\right.$ (28)
After performing the spin projection, we find the $J=1$ channel to be attractive, so there may exist a hadronic molecule of $J^{P}=1^{+}$ in the $VV$ sector.
II.3 The $cc\bar{b}\bar{b}$ and $\bar{c}\bar{c}bb$ systems
The results for the $cc\bar{b}\bar{b}$ and $\bar{c}\bar{c}bb$ systems are the same, so we only need to study the $cc\bar{b}\bar{b}$ system. In this system, the $PP$ interaction involves only one channel $B_{c}^{+}B_{c}^{+}$, whose coefficients are
$C^{t}_{PP}$ $=$ $\frac{1}{m^{2}_{J/\psi}}+\frac{1}{m^{2}_{\Upsilon}}\,,$ (29)
$C^{u}_{PP}$ $=$ $\frac{1}{m^{2}_{J/\psi}}+\frac{1}{m^{2}_{\Upsilon}}\,.$
The $VP$ interaction involves only one channel $B_{c}^{*+}B_{c}^{+}$, whose coefficient is
$C^{t}_{VP}$ $=$ $\frac{1}{m^{2}_{J/\psi}}+\frac{1}{m^{2}_{\Upsilon}}\,.$ (30)
The $VV$ interaction also involves only one channel $B_{c}^{*+}B_{c}^{*+}$, whose coefficients are
$C^{t}_{VV}$ $=$ $\frac{1}{m^{2}_{J/\psi}}+\frac{1}{m^{2}_{\Upsilon}}\,,$ (31)
$C^{u}_{VV}$ $=$ $\frac{1}{m^{2}_{J/\psi}}+\frac{1}{m^{2}_{\Upsilon}}\,,$
and the relevant contact term is
$V^{co}_{B_{c}^{*+}B_{c}^{*+}\to B_{c}^{*+}B_{c}^{*+}}(s)$ $=$ $\left\{\begin{array}[]{cc}-8g^{2}&~{}\textrm{for $J=0$},\\ 0&~{}\textrm{for $J=1$},\\ 4g^{2}&~{}\textrm{for $J=2$}.\end{array}\right.$ (35)
The above coefficients do not support the existence of hadronic molecules in the $cc\bar{b}\bar{b}$ system.
II.4 The $bb\bar{c}\bar{b}$ and $\bar{b}\bar{b}cb$ systems
The results for the $bb\bar{c}\bar{b}$ and $\bar{b}\bar{b}cb$ systems are the same, so we only need to study the $bb\bar{c}\bar{b}$ system. In this system, the $PP$ interaction involves only one channel $\eta_{b}B_{c}^{-}$, whose coefficients are
$C^{t}_{PP}$ $=$ $0\,,$ (36)
$C^{u}_{PP}$ $=$ $\lambda\frac{1}{m^{2}_{B_{c}^{*}}}\,.$
Hence, our result suggests that there do not exist hadronic molecules in the $PP$ sector.
The $VP$ interaction involves two coupled channels $\Upsilon B_{c}^{-}$ and $B_{c}^{*-}\eta_{b}$, whose coefficient is
$C_{VP}^{t}$ $=$ $\left(\begin{array}[]{c|cc}J=1&\Upsilon B_{c}^{-}&B_{c}^{*-}\eta_{b}\\ \hline\cr\Upsilon B_{c}^{-}&0&\lambda\frac{1}{m_{B_{c}^{*}}^{2}}\\ B_{c}^{*-}\eta_{b}&\lambda\frac{1}{m_{B_{c}^{*}}^{2}}&0\end{array}\right)\,.$ (40)
Diagonalizing this $2\times 2$ matrix, we obtain two decoupled channels:
$|VP^{\prime+}\rangle$ $=$ $\frac{1}{\sqrt{2}}\left(|\Upsilon B_{c}^{-}\rangle+|B_{c}^{*-}\eta_{b}\rangle\right),$ (41)
$|VP^{\prime-}\rangle$ $=$ $\frac{1}{\sqrt{2}}\left(|\Upsilon B_{c}^{-}\rangle-|B_{c}^{*-}\eta_{b}\rangle\right),$ (42)
whose coefficient is
$C_{VP}^{\prime t}$ $=$ $\left(\begin{array}[]{c|cc}J=1&VP^{\prime+}&VP^{\prime-}\\ \hline\cr VP^{\prime+}&\lambda\frac{1}{m_{B_{c}^{*}}^{2}}&0\\ VP^{\prime-}&0&-\lambda\frac{1}{m_{B_{c}^{*}}^{2}}\end{array}\right)\,.$ (46)
Hence, the interaction due to the exchange of the $B_{c}^{*}$ meson in the $VP^{\prime-}$ channel turns out to be attractive, so there may exist a hadronic molecule of $J^{P}=1^{+}$ in the $VP$ sector.
The $VV$ interaction involves only one channel $\Upsilon B_{c}^{*-}$, whose coefficients are
$C^{t}_{VV}$ $=$ $0\,,$ (47)
$C^{u}_{VV}$ $=$ $\lambda\frac{1}{m^{2}_{B_{c}^{*}}}\,,$
and the relevant contact term is
$V^{co}_{\Upsilon B_{c}^{*-}\to\Upsilon B_{c}^{*-}}(s)$ $=$ $\left\{\begin{array}[]{cc}-2g^{2}&~{}\textrm{for $J=0$},\\ 3g^{2}&~{}\textrm{for $J=1$},\\ g^{2}&~{}\textrm{for $J=2$}.\end{array}\right.$ (51)
After performing the spin projection, we find the $J=1$ channel to be attractive, so there may exist a hadronic molecule of $J^{P}=1^{+}$ in the $VV$ sector.
II.5 The $bb\bar{b}\bar{b}$ system
In the $bb\bar{b}\bar{b}$ system, the $PP$ interaction involves only one channel $\eta_{b}\eta_{b}$, the $VP$ interaction involves only one channel $\Upsilon\eta_{b}$, and the $VV$ interaction also involves only one channel $\Upsilon\Upsilon$.
Since the two vertices are both zero, $\mathcal{L}_{\eta_{b}\eta_{b}\Upsilon}=0$ and $\mathcal{L}_{\Upsilon\Upsilon\Upsilon}=0$, the interactions in this system all vanish within the extended local hidden gauge framework:
$C_{PP}^{t}$ $=$ $0\,,$
$C_{PP}^{u}$ $=$ $0\,,$
$C_{VP}^{t}$ $=$ $0\,,$ (52)
$C_{VV}^{t}$ $=$ $0\,,$
$C_{VV}^{u}$ $=$ $0\,,$
$C_{VV}^{co}$ $=$ $0\,.$
Therefore, our results do not support the existence of hadronic molecules in the $bb\bar{b}\bar{b}$ system.
III Numerical Results
In the previous section we have studied the interactions of the $cc\bar{c}\bar{c}$, $cc\bar{c}\bar{b}/\bar{c}\bar{c}cb$, $cc\bar{b}\bar{b}/\bar{c}\bar{c}bb$, $bb\bar{c}\bar{b}/\bar{b}\bar{b}cb$, and $bb\bar{b}\bar{b}$ systems. In this section we numerically study their properties. As shown in Eq. (10), the loop function $G(s)$ is regularized using the cutoff method, with the cutoff momentum $\Lambda$ describing the dynamical scale to be integrated out. Its value is quite uncertain for the exchange of fully-heavy vector mesons, and we follow Ref. Liu:2023gla to choose a broad region, $\Lambda=400\sim 1400\textrm{ MeV}$, to perform numerical analyses.
Within the extended local hidden gauge framework, the resonances are dynamically generated as poles of the scattering amplitudes $T_{PP/VP/VV}(s)$. We find the existence of eight poles that may lead to some singular structures on the invariant mass spectrum:
  • There exists a pole of $J^{P}=1^{+}$ generated in the $VP$ sector of the $cc\bar{c}\bar{b}$ system.
  • There exists a pole of $J^{P}=1^{+}$ generated in the $VV$ sector of the $cc\bar{c}\bar{b}$ system.
  • There exists a pole of $J^{P}=1^{+}$ generated in the $VP$ sector of the $bb\bar{c}\bar{b}$ system.
  • There exists a pole of $J^{P}=1^{+}$ generated in the $VV$ sector of the $bb\bar{c}\bar{b}$ system.
We summarize their positions in Table 2 with respect to the cutoff momentum $\Lambda$. Besides, there exist four charge-conjugated poles in the $\bar{c}\bar{c}cb$ and $\bar{b}\bar{b}cb$ systems.
Table 2: Pole positions with respect to the cutoff momentum $\Lambda$, in units of MeV. We only list the poles that correspond to the sub-threshold bound states.
Pole $\Lambda=400$ $\Lambda=600$ $\Lambda=800$ $\Lambda=1000$ $\Lambda=1200$ $\Lambda=1400$
$|(VP)_{cc\bar{c}\bar{b}};J^{P}=1^{+}\rangle$-- - --- - --- - --- - --- - --- -
$|(VV)_{cc\bar{c}\bar{b}};J^{P}=1^{+}\rangle$-- - --- - --- - --- - --- - --- -
$|(VP)_{bb\bar{c}\bar{b}};J^{P}=1^{+}\rangle$-- - --- - --- - $015725.3-i$ 0 $015710.8-i$ 0 $015685.2-i$ 0
$|(VV)_{bb\bar{c}\bar{b}};J^{P}=1^{+}\rangle$-- - --- - --- - --- - $015790.3-i$ 0 $015784.0-i$ 0
We find two poles in the $cc\bar{c}\bar{b}$ system: one pole in the $VP$ sector and the other in the $VV$ sector. However, both of them correspond to the virtual states when setting the cutoff momentum $\Lambda=400\sim 1400\textrm{ MeV}$, so they can only result in some threshold effects. Actually, the pole in the $VP$ sector corresponds to the sub-threshold bound state only with Λ>1550 $\Lambda>1550\textrm{ MeV}$, and the pole in the $VV$ sector corresponds to the sub-threshold bound state only with Λ>2650 $\Lambda>2650\textrm{ MeV}$. Therefore, our results do not support the existence of deeply-bound hadronic molecules in the $cc\bar{c}\bar{b}$ system.
We also find two poles in the $bb\bar{c}\bar{b}$ system: one pole in the $VP$ sector and the other in the $VV$ sector. The pole in the $VP$ sector corresponds to the sub-threshold bound state when setting Λ>850 $\Lambda>850\textrm{ MeV}$, making it possible to be identified as a hadronic molecule. This pole transfers to a virtual state and results in the threshold effect when setting Λ<850 $\Lambda<850\textrm{ MeV}$. The pole in the $VV$ sector corresponds to the sub-threshold bound state with Λ>1100 $\Lambda>1100\textrm{ MeV}$. To illustrate these two poles, we present in Fig. 2 the transition amplitudes using several different values of the cutoff momentum $\Lambda$.
(a) $T_{VP}^{2}$ with Λ=750 $\Lambda=750\textrm{ MeV}$
(b) $T_{VP}^{2}$ with Λ=850 $\Lambda=850\textrm{ MeV}$
(c) $T_{VP}^{2}$ with Λ=950 $\Lambda=950\textrm{ MeV}$
(d) $T_{VV}^{2}$ with Λ=1000 $\Lambda=1000\textrm{ MeV}$
(e) $T_{VV}^{2}$ with Λ=1100 $\Lambda=1100\textrm{ MeV}$
(f) $T_{VV}^{2}$ with Λ=1200 $\Lambda=1200\textrm{ MeV}$
Figure 2: Line shapes of the transition amplitudes $\lvert T(s)\rvert^{2}$ for the cutoff momentum $\Lambda=$ (a) 750 MeV, (b) 850 MeV, and (c) 950 MeV  in the $VP$ sector as well as $\Lambda=$ (d) 1000 MeV, (e) 1100 MeV, and (f) 1200 MeV  in the $VV|_{J=1}$ sector of the $bb\bar{c}\bar{b}$ system. The relevant thresholds are indicated by dashed lines. In the subfigures (a,b,c) the green line labeled as $T_{11}$ and the red line labeled as $T_{22}$ represent $|T_{\Upsilon B_{c}^{-}\to\Upsilon B_{c}^{-}}(s)|^{2}$ and $|T_{B_{c}^{*-}\eta_{b}\to B_{c}^{*-}\eta_{b}}(s)|^{2}$, respectively.
IV Conclusion
In this paper we systematically study the fully-heavy meson-meson interactions with the quark constituents $cc\bar{c}\bar{c}$, $cc\bar{c}\bar{b}/\bar{c}\bar{c}cb$, $cc\bar{b}\bar{b}/\bar{c}\bar{c}bb$, $bb\bar{c}\bar{b}/\bar{b}\bar{b}cb$, $bb\bar{b}\bar{b}$ through the extended local hidden gauge formalism. After solving the coupled-channel Bethe-Salpeter equation, we search for poles on both the first (physical) and second Riemann sheets. The obtained results are summarized in Table 2 with respect to the cutoff momentum $\Lambda$.
We find two poles in the $bb\bar{c}\bar{b}$ system: the pole generated in the $VP$ sector corresponds to the sub-threshold bound state when setting the cutoff momentum Λ>850 $\Lambda>850\textrm{ MeV}$, while it causes the threshold effect with Λ<850 $\Lambda<850\textrm{ MeV}$; the pole generated in the $VV$ sector corresponds to the sub-threshold bound state with Λ>1100 $\Lambda>1100\textrm{ MeV}$. These two poles are potential fully-heavy hadronic molecules, and we propose to investigate them in the $\mu^{+}\mu^{-}B_{c}^{-}$ channel at LHC. We also find two poles in the $cc\bar{c}\bar{b}$ system: the pole generated in the $VP$ sector corresponds to the sub-threshold bound state only with Λ>1550 $\Lambda>1550\textrm{ MeV}$, and the pole generated in the $VV$ sector corresponds to the sub-threshold bound state only with Λ>2650 $\Lambda>2650\textrm{ MeV}$. Besides, we find four charge-conjugated poles in the $\bar{c}\bar{c}cb$ and $\bar{b}\bar{b}cb$ systems. However, our results do not support the existence of hadronic molecules in the $cc\bar{c}\bar{c}$, $cc\bar{b}\bar{b}/\bar{c}\bar{c}bb$, and $bb\bar{b}\bar{b}$ systems.
Besides, the $cb\bar{c}\bar{b}$ system has already been investigated in our previous study Liu:2023gla , where we found the existence of the fully-heavy hadronic molecules $|B_{c}^{+}B_{c}^{-};J^{PC}=0^{++}\rangle$, $|B_{c}^{*+}B_{c}^{-}-c.c.;J^{PC}=1^{+-}\rangle$, and $|B_{c}^{*+}B_{c}^{*-};J^{PC}=2^{++}\rangle$ as well as the possible existence of $|B_{c}^{*+}B_{c}^{-}+c.c.;J^{PC}=1^{++}\rangle$. We further proposed in Ref. Liu:2023gla that a lower-mass fully-heavy meson may be able to bind two higher-mass fully-heavy hadrons. The results obtained in the present study are consistent with this proposal: the exchanged mesons of the $bb\bar{c}\bar{b}$ and $cc\bar{c}\bar{b}$ systems are both the $B_{c}^{*}$ meson, but the larger mass of the $bb\bar{c}\bar{b}$ system facilitates the formation of bound states. It is a topic of considerable interest whether the interaction of the heavy meson exchange is strong enough to form hadronic molecules. This question serves as a crucial test for the extensively investigated interaction of the light meson exchange. Therefore, the present study as well as Ref. Liu:2023gla , both of which concentrate on the interaction of the fully-heavy meson exchange, are of particular interest.
Acknowledgments
This project is supported by the National Natural Science Foundation of China under Grant No. 12075019, the Jiangsu Provincial Double-Innovation Program under Grant No. JSSCRC2021488, and the Fundamental Research Funds for the Central Universities.